OpenStudy (anonymous):

The interest rate r required to increase your investment p to the amount a in t years is found by r=(a/p)1/t−1 . Find the interest rate r for p = 8100, a = 10,000, and t = 2. Round to the nearest hundredth.

OpenStudy (anonymous):

For medal and fan (:

OpenStudy (anonymous):

goo tkhunny lol

OpenStudy (tkhunny):

Let's see if we can find that formula... \(a = p\cdot (1+r)^t\) \(\dfrac{a}{p} = (1+r)^t\) \(\left(\dfrac{a}{p}\right)^{1/t} = 1+r\) \(r = \left(\dfrac{a}{p}\right)^{1/t} - 1\) Ah! That makes SO MUCH more sense. Please find a way to write clearly. Learning just a little LaTeX would go a very long way.

OpenStudy (anonymous):

not exactly. :/ thanks for the help

OpenStudy (tkhunny):

Can we now substitute? p = 8100 a = 10,000 t = 2